OpenStudy (anonymous):

if m<4 = (3x)o and m<8 = (x + 40)o, what is the measure of <4?

OpenStudy (anonymous):

You need to give more info BUT I WILL ASSUME <4 = <8 3x = x + 40 2x = 40 x = 20 m<4 = 3(20) = (60) o